### v1

@@ III-B Semantic and Instance Segmentation
$\min_{\mathcal{U}}\;$ $\sum_{u\in\mathcal{U}}\sum_{i\in u}\big\|\mathbf{f}_{i}-\boldsymbol{\mu}_{u}\big\|_{1}+\eta\!\!\sum_{(i,j)\in\mathcal{E}}\!\!\omega_{ij}\,\mathbb{1}\!\left[\mathcal{U}(i)\neq\mathcal{U}(j)\right]$
$\omega_{ij}\;$ $=\;\exp\!\left(-\frac{\|\mathbf{f}_{i}-\mathbf{f}_{j}\|_{2}^{2}}{\sigma_{f}^{2}}\right).$ (5)
where $\mathbf{f}_{i}=[L,P,S,\|\mathbf{n}_{i}\!\cdot\!\mathbf{e}_{z}\|,z_{g,i}]^{\top}$, $\boldsymbol{\mu}_{u}$ is the median feature of segment $u$, $\eta$ is the boundary weight, and $\sigma_{f}$ is set by median feature distance.
@@
Figure 5 illustrates typical outputs for the four semantic categories used in this work. Instances are produced per class by class-specific connectivity and geometry constraints. Ground instances are not required and remain a single support layer for later height and footprint queries. Building instances are formed by agglomerating coplanar façade patches through shared boundary length >1.0>1.0 m and inter-patch normal deviation $12^{\circ}$, then merging façade sets with overlapping horizontal OBB footprints (IoU>0.3>0.3) to form a building volume; small roof patches attached to the same footprint are merged if their normals are within $25^{\circ}$ of horizontal. Vegetation instances are extracted by Euclidean clustering with a distance threshold $d_{c}=\max(0.8\,\text{m},\,1.8/\sqrt[3]{\rho})$ and a minimum of $200$ points; touching crowns are separated by a marker-based opening in height slices of $1$ m. Mobile instances are obtained with DBSCAN in $(x,y)$ with $\varepsilon=0.7$ m and $\text{minPts}=30$, refined by removing attachments to the ground boundary via a $0.15$ m erosion on the ground mask.
The LOLocation table from Sec. III-A is then made class-consistent by constraining correspondences to identical semantic labels and re-solving the assignment on each class-induced bipartite graph. Let $\mathcal{O}^{23}_{c}$ and $\mathcal{O}^{25}_{c}$ be the instance sets of class $c$ at the two epochs, and let $\delta(o_{i},o_{j})$ denote the geometric cost in Eq. (4). The class-constrained association solves
$\min_{\mathbf{X}\in\{0,1\}^{|\mathcal{O}^{23}_{c}|\times|\mathcal{O}^{25}_{c}|}}\;$ $\sum_{i\in\mathcal{O}^{23}_{c}}\sum_{j\in\mathcal{O}^{25}_{c}}\!\!\big(\delta(o_{i},o_{j})+\kappa\,\phi(o_{i},o_{j})\big)\,X_{ij}$
$+\lambda\!\!\sum_{i\in\mathcal{O}^{23}_{c}}\!\!\big(1-\sum_{j}X_{ij}\big)+\lambda\!\!\sum_{j\in\mathcal{O}^{25}_{c}}\!\!\big(1-\sum_{i}X_{ij}\big).$ (6)
where $\phi$ penalizes LoD-unsafe overlaps by $\phi(o_{i},o_{j})=\max\!\big(0,\,\mathrm{LoD}_{95}(o_{i},o_{j})-d_{\perp}(o_{i},o_{j})\big)$ with $d_{\perp}$ the median normal-direction separation over the OBB intersection, $\kappa{=}2.0$, and $\lambda$ the split/merge slack (set to $1.0$). Problem (6) is solved by a Hungarian solver on an augmented cost matrix with dummies; accepted pairs satisfy in addition OBB IoU>0.05>0.05 and centroid distance <2<2 m. This step preserves the geometric matches while enforcing semantic consistency and prepares per-instance, per-class correspondences required by the change analysis. For Subiaco, parameters are fixed as follows: $(\eta,\sigma_{f})=(0.8,\,\text{median}\,\|\mathbf{f}_{i}-\mathbf{f}_{j}\|_{2})$ in (5); ground residual $0.10$ m, building plane residual $0.08$ m and area $6$ m2; vegetation thresholds $\tau_{S}=0.35$ and crown split slice $1$ m; DBSCAN $(\varepsilon,\text{minPts})=(0.7\,\text{m},30)$; assignment $(\kappa,\lambda)=(2.0,1.0)$ with LoD gating from Eq. (3).
@@ III-C Change Analysis
For each class $c\in\{{ground},{building},{vegetation},{mobile}\}$ and each associated pair $(o_{i}^{23},o_{j}^{25})$, we compute overlap, displacement, height, volume, and shape statistics at a fixed voxel size of $0.5$ m and derive one of five labels $\{{Added},{Removed},{Increased},{Decreased},{Unchanged}\}$. All measurements are gated by the local level-of-detection $\mathrm{LoD}_{95}$ from Eq. (3) to suppress pseudo-changes.
(a) 2025 fused and height–normalized 3D LiDAR-GNSS point cloud map [38]
@@
For overlapping geometry, we approximate the normal-direction displacement by sampling the intersection of the two oriented bounding boxes and projecting the local inter-epoch offset onto the average normal. Let $\Omega(o_{i},o_{j})$ be the set of voxel centers in the OBB intersection, $p_{23}(x)$ and $p_{25}(x)$ the nearest points to $x$ in $\mathcal{P}^{23}$ and $\mathcal{P}^{25}$, and $\bar{n}(x)$ the unit average of their estimated normals. The signed normal displacement is
$D_{\perp}(o_{i},o_{j})=\operatorname{median}_{x\in\Omega(o_{i},o_{j})}\;\bar{n}(x)^{\top}\!\left(p_{25}(x)-p_{23}(x)\right),$ (7)
and we declare it informative only if $|D_{\perp}(o_{i},o_{j})|>\theta_{\text{LoD}}\cdot\overline{\mathrm{LoD}}_{95}(o_{i},o_{j})$ with $\theta_{\text{LoD}}=1.2$ and $\overline{\mathrm{LoD}}_{95}$ the median $\mathrm{LoD}_{95}$ over $\Omega(o_{i},o_{j})$. We compute 3D occupancy IoU on the union OBB at $0.5$ m, centroid shift $\Delta\mathbf{c}=\mathbf{c}_{j}-\mathbf{c}_{i}$, height change $\Delta h=h^{25}_{95}-h^{23}_{95}$, volume change $\Delta V=V^{25}-V^{23}$ where $V$ is the occupied-voxel volume, and a chi-square distance between height histograms in the intersection OBB. All scalars are reported together with the per-class decision.
Decisions are class-specific and use fixed thresholds for Subiaco. For building, a pair is labeled Increased if $\mathrm{IoU}_{3\mathrm{D}}>0.10$ and either $|\Delta h|>0.50$ m with $\Delta h>0$ and $|\Delta h|>\theta_{\text{LoD}}\cdot\overline{\mathrm{LoD}}_{95}$, or $\Delta V/V^{23}>0.10$; it is labeled Decreased if $\mathrm{IoU}_{3\mathrm{D}}>0.10$ and $\Delta V/V^{23}<-0.10$ or $\Delta h<-0.50$ m with the same LoD gate. A pair is Unchanged if $\mathrm{IoU}_{3\mathrm{D}}>0.20$, $\|\Delta\mathbf{c}\|<1.5$ m, $|\Delta h|\leq 0.50$ m, and $|\Delta V|/V^{23}\leq 0.10$, or if all informative statistics remain below their LoD. If $\mathrm{IoU}_{3\mathrm{D}}\leq 0.10$ but a building instance exists only at 2025, it is Added; if it exists only at 2023, it is Removed. For vegetation, we keep the same logic with $(\mathrm{IoU}_{3\mathrm{D}}>0.08,\ \|\Delta\mathbf{c}\|<2.0$ m)) and thresholds $(|\Delta h|>0.30$ m, $|\Delta V|/V^{23}>0.15)$; the histogram distance is used to break ties in favor of Increased when vertical mass shifts upward. For ground, changes are evaluated on a $2$ m raster: a tile is Increased if the median ground elevation difference exceeds +0.15+0.15 m over a contiguous area >25>25 m2; it is Decreased if the difference is below 0.15-0.15 m with the same area constraint; otherwise Unchanged. Ground Added/Removed does not apply. For mobile, instances unmatched across epochs are labeled Added/Removed; matched instances are Unchanged if $\mathrm{IoU}_{3\mathrm{D}}>0.20$ and $\|\Delta\mathbf{c}\|<2.0$ m, otherwise they are discarded from city-scale reporting since persistent mobiles are rare in multi-year intervals.
Figure 7: Qualitative change detection on three representative Subiaco blocks. Each image presents the 2023 and 2025 projections, height maps, semantic and instance views, and the object-level change map with the legend embedded. From top to bottom: Blocks 6, 7, and 8.
To aggregate heterogeneous evidence into a confidence score attached to each decision, we standardize the informative statistics by their uncertainty and combine them through a logistic map. Let $z_{h}=\frac{|\Delta h|}{\sqrt{\overline{\mathrm{LoD}}_{95}^{\,2}+\sigma_{h}^{2}}}$, $z_{v}=\frac{|\Delta V|}{\sqrt{(V^{23}\cdot\sigma_{v})^{2}+\epsilon}}$, $z_{o}=1-\mathrm{IoU}_{3\mathrm{D}}$, and $z_{c}=\frac{\max(0,\|\Delta\mathbf{c}\|-\tau_{c})}{\tau_{c}}$ with $(\tau_{c}^{\text{bld}},\tau_{c}^{\text{veg}})=(1.5\,\text{m},2.0\,\text{m})$. The confidence for the selected label $y$ on pair $(o_{i},o_{j})$ is
$s_{y}(o_{i},o_{j})$ $=(1+\exp[-(w_{h}z_{h}+w_{v}z_{v}+w_{o}z_{o}$
$\qquad\quad+\,w_{c}z_{c}+\,w_{\perp}\,\tfrac{|D_{\perp}(o_{i},o_{j})|}{\overline{\mathrm{LoD}}_{95}})])^{-1}.$ (8)
with class-dependent weights fixed for Subiaco as $(w_{h},w_{v},w_{o},w_{c},w_{\perp})=(0.35,0.30,0.20,0.10,0.05)$ for buildings and $(0.30,0.35,0.15,0.10,0.10)$ for vegetation; for ground we use $(w_{h},w_{v},w_{o},w_{c},w_{\perp})=(0.50,0.00,0.25,0.00,0.25)$ with $\Delta h$ interpreted as median elevation difference on the tile. The score $s_{y}$ is reported jointly with the label.
@@ IV Experiments
This section evaluates the approach on the Subiaco map pair under a consistent protocol. Qualitative results on representative blocks illustrate typical urban patterns, and quantitative results report accuracy, macro F1, and macro IoU together with per class IoU. Comparisons with classical and point based baselines assess effectiveness across object categories and ground works, using fixed parameters and tiled processing to reflect the intended operational setting.
IV-A Subiaco 2025 3D Map Construction Result
@@
Figure 6 compares the 2025 reconstruction with the 2023 map after registration into the 2025 frame. The 2025 map shows continuous curb lines, coherent façades across contiguous blocks and closed intersections with consistent approach geometry. Roof ridges and small roof appendages are preserved. Under canopy the ground surface remains connected and tree crowns form complete outer envelopes. The 2023 overlay presents local shear near loop junctions, duplicated façade strips from drift and lateral offsets at intersections. The 2025 reconstruction is used as the geometric reference in all subsequent analyses.
IV-B Qualitative Change Detection on Representative Blocks
@@
To enable city–scale processing and visualization, both epochs are partitioned into 52 square blocks of $80\times 80$ m with a $10$ m overlap to avoid boundary artifacts during association and to keep each tile under memory limits at 0.25 m voxels. All tiles are registered into the 2025 frame, and change labels are produced per object with the LoD gate and the class rules defined in the methodology. Results are visualized with five panels per block: raw projection, height map, semantic labels, instance labels, and the change map that encodes Added in green, Removed in red, Increased in blue, Decreased in orange, and Unchanged in gray. Blocks 6, 7, and 8 are reported because they cover the typical patterns observed in Subiaco, namely building redevelopment around courtyards and street corners, vegetation growth and pruning under canopy, and roadway resurfacing at multi-lane intersections with parking bays. They also include strong occlusion, long façades, and tree rows, which are failure modes for point-wise differencing but are handled by object-level reasoning.
Figure 7 shows qualitative outcomes for the three blocks. In Block 6 the method flags a large vegetation cluster as Removed at the center, multiple crowns along the verges as Increased, and narrow ground strips along the carriageway as Decreased consistent with resurfacing, while building volumes remain Unchanged. In Block 7 new built volumes inside the inner lot are labeled Added with height Increased at the northern frontage, and several shrubs along the southern footpath are Removed; ground in the adjacent parking area is Decreased and the surrounding façades keep high overlap. In Block 8 tree rows along a boulevard produce alternating Added and Removed crowns, a central median is reprofiled with contiguous Increased ground, and two small structures at the edge are Removed. Across the three tiles thin façade duplicates and boundary splits are suppressed by the overlap and by class-consistent assignment, and small ephemeral objects are ignored by the instance size and LoD thresholds. These examples illustrate the intended operating regime before the quantitative evaluation.
@@
TABLE I: Comparison with representative methods on the Subiaco dataset using identical metrics. Accuracy is micro accuracy, mF1 is macro F1 across the five change labels, and mIoU is macro IoU. Right-hand columns report per-class IoU.
@@
Object–based (ours) $95.2$ $90.4$ $82.6$ $86.9$ $80.8$ $76.6$ $74.8$ $93.8$
TABLE II: Performance on 15 representative Subiaco blocks (out of 52). Metrics are accuracy (micro accuracy), mF1 (macro F1 across five change labels), and mIoU (macro IoU). Right-hand columns report per-class IoU. All values are percentages. The last row aggregates all instances across the 15 blocks.
Block ACC mF1 mIoU Per Class IoU (%)
Added Removed Increased Decreased Unchanged
B06 95.1 90.3 82.4 88.4 81.8 75.3 74.0 92.5
B07 95.4 90.7 83.1 87.6 81.9 74.5 77.7 93.8
B08 94.8 89.9 81.6 86.7 81.8 74.0 70.2 95.3
B09 95.7 91.3 84.0 85.9 82.7 78.5 78.2 94.9
B10 94.1 88.6 79.6 87.2 82.7 75.9 60.3 92.1
B11 96.0 91.6 84.8 85.1 79.1 76.2 90.0 93.7
B12 94.5 89.4 80.7 85.9 79.1 76.3 69.2 92.9
B13 95.6 91.0 83.5 88.4 80.8 77.2 79.1 92.1
B14 95.2 90.8 83.0 88.4 78.6 75.7 76.3 96.0
B15 94.3 88.9 80.4 88.7 80.1 78.2 60.0 94.8
B16 94.9 90.1 81.9 87.4 82.4 78.2 68.5 93.2
B17 95.8 91.5 84.6 85.1 79.2 78.0 86.5 94.4
B18 95.3 90.7 82.9 87.2 81.5 77.4 75.7 92.7
B19 94.7 89.7 81.2 87.0 81.9 76.6 66.9 93.8
B20 95.9 91.4 84.8 85.1 78.2 77.5 89.4 94.0
all 95.2 90.4 82.6 86.9 80.8 76.6 74.8 93.8
IV-C Quantitative Evaluation on City Blocks and Method Comparison
We evaluate object–level change prediction on the Subiaco map pair using the five labels {Added, Removed, Increased, Decreased, Unchanged}. Three metrics are used consistently: accuracy is the micro overall accuracy, mF1 is the macro–averaged F1 across the five labels, and mIoU is the macro–averaged Intersection over Union. Per–class IoU is reported to expose failure modes. Fifteen representative blocks are selected from the 52 tiled blocks to cover redevelopments, vegetation dynamics and ground works while keeping the class distribution comparable to the full set. Across the 15 blocks in Table II, accuracy ranges from 94.1% to 96.0%, mF1 from 88.6% to 91.6%, and mIoU from 79.6% to 84.8%. IoU for Unchanged is above 92% on all blocks, indicating stable assignment of non–events. The lowest per–class scores appear on Decreased where ground reprofiling is narrow and partially overlaps with curbs. The last row aggregates all instances across the 15 blocks and gives the overall performance of our pipeline: 95.2% accuracy, 90.4% mF1 and 82.6% mIoU with per–class IoUs of 86.9% for Added, 80.8% for Removed, 76.6% for Increased, 74.8% for Decreased, and 93.8% for Unchanged.
We compare the pipeline with representative baselines and point-based models under a unified protocol. Table I shows that classical DSM and RF variants remain below 60% mIoU. KPConv variants raise mIoU to the low 80s but are sensitive to thin ground changes. Using the aggregated metrics from Table II, our object-based method achieves 95.2% accuracy, 90.4% mF1, and 82.6% mIoU, surpassing all compared methods on all three summary metrics. Relative to Triplet KPConv, the improvements are 0.2 percentage points in accuracy, 0.2 percentage points in mF1, and 0.8 percentage points in mIoU. The largest classwise gain is observed for Decreased, where the IoU reaches 74.8%, exceeding the baseline by 7.6 percentage points and reducing false positives on narrow resurfaced strips while preserving competitive performance on the remaining classes.

### v2

@@ III-B Semantic and Instance Segmentation
$\min_{\mathcal{U}}\;$ $\sum_{u\in\mathcal{U}}\sum_{i\in u}\bigl\|\mathbf{f}_{i}-\boldsymbol{\mu}_{u}\bigr\|_{1}+\eta\!\!\sum_{(i,j)\in\mathcal{E}}\!\!\omega_{ij}\,\mathbb{1}\!\left[\mathcal{U}(i)\neq\mathcal{U}(j)\right],$
$\text{where}\quad\omega_{ij}\;$ $=\exp\!\Bigl(-\|\mathbf{f}_{i}-\mathbf{f}_{j}\|_{2}^{2}/\sigma_{f}^{2}\Bigr).$ (5)
$\mathbf{f}_{i}=[L,P,S,\|\mathbf{n}_{i}\!\cdot\!\mathbf{e}_{z}\|,z_{g,i}]^{\top}$, $\boldsymbol{\mu}_{u}$ is the median feature of segment $u$, $\eta$ is the boundary weight, and $\sigma_{f}$ is set by median feature distance.
@@
Fig 6 illustrates typical outputs for the four semantic categories used in this work. Instances are produced per class by class-specific connectivity and geometry constraints. Ground instances are not required and remain a single support layer for later height and footprint queries. Building instances are formed by agglomerating coplanar facade patches through shared boundary length >1.0>1.0 m and inter-patch normal deviation $12^{\circ}$, then merging facade sets with overlapping horizontal OBB footprints (IoU>0.3>0.3) to form a building volume; small roof patches attached to the same footprint are merged if their normals are within $25^{\circ}$ of horizontal. Vegetation instances are extracted by Euclidean clustering with a distance threshold $d_{c}=\max(0.8\,\text{m},\,1.8/\sqrt[3]{\rho})$ and a minimum of $200$ points; touching crowns are separated by a marker-based opening in height slices of $1$ m. Mobile instances are obtained with DBSCAN in $(x,y)$ with $\varepsilon=0.7$ m and $\text{minPts}=30$, refined by removing attachments to the ground boundary via a $0.15$ m erosion on the ground mask.
The level object location table is made class consistent by restricting associations to instances with the same semantic label and solving a regularized assignment problem per class. For class $c$, let $\mathcal{O}_{c,2023}$ and $\mathcal{O}_{c,2025}$ denote the instance sets at the two epochs. The binary matrix $\mathbf{X}\in\{0,1\}^{|\mathcal{O}_{c,2023}|\times|\mathcal{O}_{c,2025}|}$ encodes matches, and $u_{i}(\mathbf{X})=1-\sum_{j\in\mathcal{O}_{c,2025}}X_{ij}$ and $v_{j}(\mathbf{X})=1-\sum_{i\in\mathcal{O}_{c,2023}}X_{ij}$ measure unmatched instances on each side. The class-constrained association is obtained by
$\min_{\mathbf{X}}$ $\sum_{i\in\mathcal{O}_{c,2023}}\sum_{j\in\mathcal{O}_{c,2025}}\big(\delta(o_{i},o_{j})+\phi(o_{i},o_{j})\big)X_{ij}$ (6)
$\;+\lambda\sum_{i\in\mathcal{O}_{c,2023}}u_{i}(\mathbf{X})+\lambda\sum_{j\in\mathcal{O}_{c,2025}}v_{j}(\mathbf{X})$
s.t. $\mathbf{X}\in\{0,1\}^{|\mathcal{O}_{c,2023}|\times|\mathcal{O}_{c,2025}|}.$
Here $\phi$ penalizes LoD-unsafe overlaps through $\phi(o_{i},o_{j})=\max\big(0,\,\mathrm{LoD}_{95}(o_{i},o_{j})-d_{\perp}(o_{i},o_{j})\big)$, where $d_{\perp}(o_{i},o_{j})$ is the median normal-direction separation over the intersection of the two oriented bounding boxes. The weights are set to $\kappa=2.0$ and $\lambda=1.0$. Problem (6) is solved with the Hungarian algorithm on an augmented cost matrix with dummy rows and columns, and a match is accepted only if the 3D OBB IoU exceeds $0.05$ and the centroid distance is below $2$ m. This produces per-instance, per-class correspondences that respect both semantics and geometry and are passed to the change analysis layer.
For all experiments on Subiaco the remaining parameters are fixed as follows. In (5) we use $(\eta,\sigma_{f})=(0.8,\,\text{median}\,\|\mathbf{f}_{i}-\mathbf{f}_{j}\|_{2})$. The ground plane residual is $0.10$ m. The building plane residual is $0.08$ m with minimum planar area $6$ m2. Vegetation uses $\tau_{S}=0.35$ and a crown split slice of $1$ m. DBSCAN operates with $(\varepsilon,\text{minPts})=(0.7\,\text{m},30)$. The assignment parameters are $(\kappa,\lambda)=(2.0,1.0)$ with LoD gating given by (3).
@@ III-C Change Analysis
For each class $c\in\{{ground},{building},{vegetation},{mobile}\}$ and each associated pair $(o_{i,2023},o_{j,2025})$, we compute overlap, displacement, height, volume, and shape statistics at a fixed voxel size of $0.5$ m and derive one of five labels $\{{Added},{Removed},{Increased},{Decreased},\\ {Unchanged}\}$. The complete dataset is manually labeled into these five classes for evaluation of our method and comparison to other techniques. We also release these labels along with the Subiaco 2025 map dataset. All measurements are gated by the local level-of-detection $\mathrm{LoD}_{95}$ from Eq. (3) to suppress pseudo-changes.
For overlapping geometry, we approximate the normal-direction displacement by sampling the intersection of the two oriented bounding boxes and projecting the local inter-epoch offset onto the average normal. Let $\Omega(o_{i},o_{j})$ be the set of voxel centers in the OBB intersection, $p_{2023}(x)$ and $p_{2025}(x)$ the nearest points to $x$ in $\mathcal{P}_{2023}$ and $\mathcal{P}_{2025}$, and $\bar{n}(x)$ the unit average of their estimated normals. The signed normal displacement is
$\vskip-2.84526ptD_{\perp}(o_{i},o_{j})=\operatorname*{median}_{x\in\Omega(o_{i},o_{j})}\;\bar{n}(x)^{\top}\!\left(p_{2025}(x)-p_{2023}(x)\right).$ (7)
and we declare it informative only if $|D_{\perp}(o_{i},o_{j})|>\theta_{\text{LoD}}\cdot\overline{\mathrm{LoD}}_{95}(o_{i},o_{j})$ with $\theta_{\text{LoD}}=1.2$ and $\overline{\mathrm{LoD}}_{95}$ the median $\mathrm{LoD}_{95}$ over $\Omega(o_{i},o_{j})$. We compute 3D occupancy IoU on the union OBB at $0.5$ m, centroid shift $\Delta\mathbf{c}=\mathbf{c}_{j}-\mathbf{c}_{i}$, height change $\Delta h=h^{2025}_{95}-h^{2023}_{95}$, volume change $\Delta V=V^{2025}-V^{2023}$ where $V$ is the occupied-voxel volume, and a chi-square distance between height histograms in the intersection OBB. All scalars are reported together with the per-class decision.
Decisions are class-specific and use fixed thresholds for Subiaco. For building, a pair is labeled Increased if $\mathrm{IoU}_{3\mathrm{D}}>0.10$ and either $|\Delta h|>0.50$ m with $\Delta h>0$ and $|\Delta h|>\theta_{\text{LoD}}\cdot\overline{\mathrm{LoD}}_{95}$, or $\Delta V/V^{2023}>0.10$; it is labeled Decreased if $\mathrm{IoU}_{3\mathrm{D}}>0.10$ and $\Delta V/V^{2023}<-0.10$ or $\Delta h<-0.50$ m with the same LoD gate. A pair is Unchanged if $\mathrm{IoU}_{3\mathrm{D}}>0.20$, $\|\Delta\mathbf{c}\|<1.5$ m, $|\Delta h|\leq 0.50$ m, and $|\Delta V|/V^{2023}\leq 0.10$, or if all informative statistics remain below their LoD. If $\mathrm{IoU}_{3\mathrm{D}}\leq 0.10$ but a building instance exists only at 2025, it is Added; if it exists only at 2023, it is Removed. For vegetation, we keep the same logic with $(\mathrm{IoU}_{3\mathrm{D}}>0.08,\ \|\Delta\mathbf{c}\|<2.0$ m)) and thresholds $(|\Delta h|>0.30$ m, $|\Delta V|/V^{2023}>0.15)$; the histogram distance is used to break ties in favor of Increased when vertical mass shifts upward. For ground, changes are evaluated on a $2$ m raster: a tile is Increased if the median ground elevation difference exceeds +0.15+0.15 m over a contiguous area >25>25 m2; it is Decreased if the difference is below 0.15-0.15 m with the same area constraint; otherwise Unchanged. Ground Added/Removed does not apply. For mobile, instances unmatched across epochs are labeled Added/Removed; matched instances are Unchanged if $\mathrm{IoU}_{3\mathrm{D}}>0.20$ and $\|\Delta\mathbf{c}\|<2.0$ m, otherwise they are discarded from city-scale reporting since persistent mobiles are rare in multi-year intervals.
To aggregate heterogeneous evidence into a confidence score attached to each decision, we standardize the informative statistics by their uncertainty and combine them through a logistic map. Let $z_{h}=\frac{|\Delta h|}{\sqrt{\overline{\mathrm{LoD}}_{95}^{\,2}+\sigma_{h}^{2}}}$, $z_{v}=\frac{|\Delta V|}{\sqrt{(V^{2023}\cdot\sigma_{v})^{2}+\epsilon}}$, $z_{o}=1-\mathrm{IoU}_{3\mathrm{D}}$, and $z_{c}=\frac{\max(0,\|\Delta\mathbf{c}\|-\tau_{c})}{\tau_{c}}$ with $(\tau_{c}^{\text{bld}},\tau_{c}^{\text{veg}})=(1.5\,\text{m},2.0\,\text{m})$. The confidence for the selected label $y$ on pair $(o_{i},o_{j})$ is
$\begin{split}s_{y}(o_{i},o_{j})=\Bigl(1+\exp\!\bigl[-\bigl(w_{h}z_{h}+w_{v}z_{v}+w_{o}z_{o}+w_{c}z_{c}\\ \qquad\qquad\quad+\,w_{\perp}\,\tfrac{|D_{\perp}(o_{i},o_{j})|}{\overline{\mathrm{LoD}}_{95}}\bigr)\bigr]\Bigr)^{-1}.\end{split}$ (8)
with class-dependent weights fixed for Subiaco as $(w_{h},w_{v},w_{o},w_{c},w_{\perp})=(0.35,0.30,0.20,\\ 0.10,0.05)$ for buildings and $(0.30,0.35,0.15,0.10,0.10)$ for vegetation; for ground we use $(w_{h},w_{v},w_{o},w_{c},w_{\perp})=(0.50,0.00,0.25,0.00,0.25)$ with $\Delta h$ interpreted as median elevation difference on the tile. The score $s_{y}$ is reported jointly with the label.
(a) 2025 fused and height–normalized 3D LiDAR-GNSS map [39]
@@ IV Experiments
We evaluate our method on the proposed Subiaco map dataset and compare it to existing techniques using paired maps from 2023 and 2025 under a consistent protocol. Qualitative results on representative blocks illustrate typical urban patterns, and quantitative results report accuracy, macro F1, and macro IoU together with per class IoU. Comparisons with classical and point based baselines assess effectiveness across object categories and ground works, using fixed parameters and tiled processing to reflect the intended operational setting.
IV-A Subiaco 2025 3D Map Construction Results
@@
Fig 7 compares the 2025 reconstruction with the 2023 map. The 2025 map shows continuous curb lines, coherent facades across contiguous blocks and closed intersections with consistent approach geometry. Roof ridges and small roof appendages are preserved. Under canopy the ground surface remains connected and tree crowns form complete outer envelopes. The 2023 overlay presents local shear near loop junctions, duplicated facade strips from drift and lateral offsets at intersections. The 2025 reconstruction is used as the geometric reference in all subsequent analyses.
IV-B Qualitative Change Detection on Representative Block Samples
@@
To enable city–scale processing and visualization, both epochs are partitioned into 52 square blocks of $80\times 80$ m with a $10$ m overlap to avoid boundary artifacts during association and to keep each tile under memory limits at 0.25 m voxels. All tiles are registered into the 2025 frame, and change labels are produced per object with the LoD gate and the class rules defined in the methodology. Results are visualized with five panels per block: raw projection, height map, semantic labels, instance labels, and the change map that encodes Added in green, Removed in red, Increased in blue, Decreased in orange, and Unchanged in gray. Block samples from loop 6, 7, and 8 are reported because they cover the typical patterns observed in Subiaco, namely building redevelopment around courtyards and street corners, vegetation growth and pruning under canopy, and roadway resurfacing at multi-lane intersections with parking bays. They also include strong occlusion, long facades, and tree rows, which are failure modes for point-wise differencing but are handled by object-level reasoning.
@@
TABLE I: Performance on the Subiaco dataset across 20 loops. Metrics are accuracy (micro accuracy), mF1 (macro F1 across five change labels), and mIoU (macro IoU). Right-hand columns report per-class IoU. All values are percentages.
Loop ACC mF1 mIoU Per Class IoU (%)
Added Removed Increased Decreased Unchanged
01 95.7 92.2 84.1 89.4 83.0 78.1 78.8 95.0
02 96.0 92.0 83.8 89.1 82.7 77.8 78.1 96.9
03 95.8 91.9 83.6 88.7 82.5 77.6 77.4 96.8
04 96.2 92.3 84.3 89.6 83.2 78.4 79.2 95.1
05 94.5 91.8 83.5 89.0 82.4 77.3 77.0 97.1
06 95.1 90.3 82.4 88.4 81.8 75.3 74.0 92.5
07 95.4 90.7 83.1 87.6 81.9 74.5 77.7 93.8
08 94.8 89.9 81.6 86.7 81.8 74.0 70.2 95.3
09 95.7 91.3 84.0 85.9 82.7 78.5 78.2 94.9
10 94.1 88.6 79.6 87.2 82.7 75.9 60.3 92.1
11 96.0 91.6 84.8 85.1 79.1 76.2 90.0 93.7
12 94.5 89.4 80.7 85.9 79.1 76.3 69.2 92.9
13 95.6 91.0 83.5 88.4 80.8 77.2 79.1 92.1
14 95.2 90.8 83.0 88.4 78.6 75.7 76.3 96.0
15 94.3 88.9 80.4 88.7 80.1 78.2 60.0 94.8
16 94.9 90.1 81.9 87.4 82.4 78.2 68.5 93.2
17 95.8 91.5 84.6 85.1 79.2 78.0 86.5 94.4
18 95.3 90.7 82.9 87.2 81.5 77.4 75.7 92.7
19 94.7 89.7 81.2 87.0 81.9 76.6 66.9 93.8
20 95.9 91.4 84.8 85.1 78.2 77.5 89.4 94.0
all 95.3 90.8 82.9 87.5 81.3 76.9 75.6 94.3
Figure 8: Qualitative change detection on three representative block samples. Each image presents the 2023 and 2025 projections, height maps, semantic and instance views, and the object-level change map with the legend embedded. From top to bottom: Samples from loop 6, 7, and 8.
Fig 8 shows qualitative outcomes for the three samples. The top one shows the method flags a large vegetation cluster as Removed at the center, multiple crowns along the verges as Increased, and narrow ground strips along the carriageway as Decreased consistent with resurfacing, while building volumes remain Unchanged. The middle visualization shows that new volumes within the inner lot are marked Added with height Increased at the northern frontage; shrubs along the southern footpath are Removed, the adjacent parking ground is Decreased, and surrounding facades retain high overlap. In the bottom section, the sample shows tree rows along the boulevard forming alternating Added and Removed crowns; the central median is reprofiled with contiguous Increased ground, and two small edge structures are Removed. Across the three tiles thin facade duplicates and boundary splits are suppressed by the overlap and by class-consistent assignment, and small ephemeral objects are ignored by the instance size and LoD thresholds. These examples illustrate the intended operating regime before the quantitative evaluation.
TABLE II: Comparison with representative methods on the Subiaco dataset using identical metrics. All methods are evaluated on the first five loops (01–05). Accuracy is micro accuracy, mF1 is macro F1 across the five change labels, and mIoU is macro IoU. Right-hand columns report per-class IoU.
@@
Object–based (ours) 95.6 92.0 83.9 89.1 82.8 77.8 78.1 96.2
IV-C Quantitative Evaluation on all 21 Loops
We first evaluate the proposed method on all twenty loops using the five change labels. Accuracy, macro F1 and macro IoU are computed per loop, and per class IoU is reported to expose systematic errors. Table I shows that accuracy remains within 94.1% to 96.2%, macro F1 within 88.6% to 92.3%, and macro IoU within 79.6% to 84.8%. The Unchanged class consistently exceeds 92% IoU, indicating stable behavior on non events. Lower scores on Decreased reflect narrow resurfacing and partial curb overlap where geometry is sparse. Aggregating all loops yields 95.3% accuracy, 90.8% macro F1 and 82.9% macro IoU, with class IoUs of 87.5% for Added, 81.3% for Removed, 76.9% for Increased, 75.6% for Decreased and 94.3% for Unchanged. These results quantify the performance of our model when trained on loops 6 to 20 and evaluated across the full set.
IV-D Comparison to Learning based Methods on the Test Set
Although our method does not require training data, for comparison to learning based methods, we partitioned the map into mutually exclusive training and test sets as shown in Fig 2. In Table II, we compare our method to learning based methods on the test partition. Classical DSM based and RF based approaches remain below 60% macro IoU. KPConv and its variants increase performance to the low 80s but remain sensitive to thin ground changes. Our object based method reaches 95.6% accuracy, 92.0% macro F1 and 83.9% macro IoU on the same test set. Relative to Triplet KPConv, the gains are 0.6 points in accuracy, 1.8 points in macro F1 and 2.1 points in macro IoU. The largest improvement appears on the Decreased category where our method achieves 78.1% IoU, outperforming the nearest competitor by 10% points. Our method reduces false detections on narrow resurfaced strips while maintaining consistent results across the remaining classes.